### v2

@@ Abstract
Time series forecasting is a critical task that provides key information for decision-making across various fields, such as economic planning, supply chain management, and medical diagnosis. After the use of traditional statistical methodologies and machine learning in the past, various fundamental deep learning architectures such as MLPs, CNNs, RNNs, and GNNs have been developed and applied to solve time series forecasting problems. However, the structural limitations caused by the inductive biases of each deep learning architecture constrained their performance. Transformer models, which excel at handling long-term dependencies, have become significant architectural components for time series forecasting. However, recent research has shown that alternatives such as simple linear layers can outperform Transformers. These findings have opened up new possibilities for using diverse architectures. In this context of exploration into various models, the architectural modeling of time series forecasting has now entered a renaissance. This survey not only provides a historical context for time series forecasting but also offers comprehensive and timely analysis of the movement toward architectural diversification. By comparing and re-examining various deep learning models, we uncover new perspectives and presents the latest trends in time series forecasting, including the emergence of hybrid models, diffusion models, Mamba models, and foundation models. By focusing on the inherent characteristics of time series data, we also address open challenges that have gained attention in time series forecasting, such as channel dependency, distribution shift, causality, and feature extraction. This survey explores vital elements that can enhance forecasting performance through diverse approaches. These contributions lead to lowering the entry barriers for newcomers to the field of time series forecasting, while also offering seasoned researchers broad perspectives, new opportunities, and deep insights.
@@ 1 Introduction
Time series forecasting (TSF) is a task that predicts future values based on sequential historical data over time [34]. It is utilized as a key decision-making tool in various fields, such as economics and finance, supply chain management, transportation, energy, weather and healthcare [35, 1, 6, 135, 37, 164]. Such applications offer various opportunities, including cost reduction, increased efficiency, and enhanced competitiveness [35]. The inherent diversity and complexity of time series data unfortunately make forecasting challenging. In addition to the apparent information, various hidden patterns make it challenging to learn temporal dependencies, and irregular values at times further complicate the problem. In multivariate problems, additional factors such as channel correlation make the task even more difficult (Section 2.1). Furthermore, time series data exhibits different characteristics depending on the domain, and the various times and environments in which the series is collected result in significantly different patterns (Section 2.3). Because of this, TSF problems exhibit limited model generalizability, requiring diverse architectures and approaches. The increasingly complicated TSF problems are presenting researchers with growing challenges, which has recently led to the active development of new methodologies and algorithms to address these issues [104].
@@
The explosive increase in the number of papers accepted at major AI and Machine Learning conferences (Fig. 2) demonstrates that TSF research is becoming increasingly important in the AI and Machine Learning fields. As various studies addressing time series forecasting problems are being actively conducted, survey papers are also being frequently published (Table 1). Over time, numerous survey papers have systematically organized the vast landscape of TSF, offering in-depth research that has provided valuable guidance and direction for researchers. However, existing survey papers still have room for improvement, particularly in addressing the inevitable increase in model diversity and the open challenges in the field.
@@
Models for TSF have undergone various stages of development over an extended period. In the past, statistical methods based on moving averages were predominantly used, which later evolved into traditional approaches such as Exponential Smoothing and ARIMA [11]. Machine learning techniques such as Tree Models [144] and Support Vector Machines (SVM) [31] have also been frequently used, but they had limitations in learning complex nonlinear patterns (Section 3.1). With the increase in available data and advancements in hardware computing power, various deep learning architectures such as MLPs [152], RNNs [74], CNNs [94], and GNNs [155] were developed, enabling the learning of more complex patterns. However, the performance of these early deep learning architectures was constrained by their intrinsic designs. To overcome these structural limitations, variants such as Long Short-Term Memory (LSTM) [71] and Temporal Convolutional Networks (TCN) [9] have been widely utilized (Section 3.2). Transformers [177], known for their ability to handle long-term dependencies, have demonstrated excellent performance in natural language processing and have been naturally extended to time series data as well. While Transformers have shown good performance in TSF and become widely popular, recent cases have shown that simple linear models can outperform Transformer models (Section 3.3). As a result, there has been a significant increase in reconsidering traditional deep learning methodologies, along with growing interest in various architectures such as foundation models, diffusion models and Mamba models (Section 4.2, 4.3, 4.4, 4.5). The Transformer model continues to improve in performance and still plays a significant role (Section 4.1). In this way, TSF has entered a renaissance of modeling, with various methodologies actively competing without being dominated by any single approach (Fig. 3). In this context, this survey offers two major strengths that set it apart from previous TSF survey papers.
@@
Second, we explore from the perspective of open challenges. Although numerous advanced architectures have resolved many issues, the core challenges in TSF continue to persist. In particular, issues such as channel correlation, distribution shifts, causality, and feature extraction (Section 5) remain significant challenges that need to be addressed (Fig. 4). This survey explores the latest methodologies aimed at addressing these challenges and provides readers with valuable insights for problem-solving. While previous surveys have provided useful perspectives on open challenges, they have not explored these issues in sufficient depth. This survey aims to bridge that gap by offering a more comprehensive analysis and proposing new solutions.
@@
    Articles Focus Broad Architecture Review Forecasting Task-Intensive Recent Work Reference
@@
    Forecast Methods for Time Series Data: A Survey · Traditional forecasting methods: statistical, ML, and DL
@@ A Review on Deep Sequential Models for Forecasting Time Series Data · Common deep sequential models
(Section 5.1)(Section 5.2)(Section 5.3)(Section 5.4)
@@
This survey covers the fundamental concepts of time series data and the problem definition of forecasting in Section 2, followed by an examination of the evolution of past methodologies in Section 3. In Section 4, it analyzes the key features of the latest models, and finally, in Section 5, it explores the open challenges in TSF and their solutions. Through this, readers will gain a broad understanding of the past and present of TSF research and acquire new ideas for future studies.
@@ 2.1 Time-Series Data
Time series data is a collection of sequential data points gathered at regular time intervals, representing a series of observations of phenomena that change over time. This temporal continuity allows for the understanding and analysis of phenomena that evolve according to time order. Each data point represents the state or value at a specific moment, and through the observation of these data points, various patterns such as long-term trends, seasonality, cyclicality, and irregularities can be recognized. These patterns provide valuable information for predicting future values or detecting changes at critical moments. By extracting and studying the meaningful information provided by time series data, it can be applied to practical applications, helping to address many challenges in various disciplines.
@@ 2.2 Time-Series Forecasting
Univariate forecasting refers to making predictions using only one variable. For example, predicting the next day’s temperature based solely on past temperature data from a weather station is a univariate forecast. The advantage of univariate forecasting is that the models are simple and computationally efficient. Since the data consists of a single variable, the model is relatively straightforward and easy to understand, making data collection and management easier. However, it may only utilize limited information as it cannot account for important external factors or interactions between different variables:
$\hat{y}_{t+1}=f(y_{t},y_{t-1},\ldots,y_{t-p})$ (1)
where $\hat{y}_{t+1}$ represents the predicted value at time $t+1$. The terms $y_{t},y_{t-1},\ldots,y_{t-p}$ are the past $p$ observations, which include values at time $t$, $t-1$, continuing down to $t-p$. This approach uses historical data to forecast future values, relying on patterns within the observed variable itself.
@@
To incorporate other influencing variables, we can extend the model to a multivariate approach, as shown in the following equation:
$\hat{y}_{t+1}=f(y_{t},y_{t-1},\ldots,y_{t-p},x_{t},x_{t-1},\ldots,x_{t-p})$ (2)
where $\hat{y}_{t+1}$ is the predicted value at time $t+1$. The terms $y_{t},y_{t-1},\ldots,y_{t-p}$ represent the past $p$ observations of the target variable, capturing its historical patterns. Additionally, $x_{t},x_{t-1},\ldots,x_{t-p}$ are the past $p$ observations of other related variables, which provide additional context and information that may influence the prediction. This multivariate approach allows for more comprehensive modeling by considering both the target variable’s history and the effects of other relevant factors. Fig. 7 illustrates how a multivariate forecasting model uses a past lookback window to predict future intervals.
@@ 2.3 Time Series Forecasting Datasets
Time series data exhibit a greater variety of domain-specific characteristics compared to Natural Language Processing (NLP). These differences in dataset characteristics make time series forecasting problems more diverse and complex, reducing the generalizability of time series forecasting models. Consequently, developing foundation models for TSF is more challenging compared to NLP. Time series data lack the well-designed vocabulary and grammar found in NLP, and obtaining vast amounts of data is more difficult than in fields like Computer Vision(CV) or NLP. In the CV field, there are general benchmark datasets such as MNIST and ImageNet, while in NLP, there are datasets like GLUE(General Language Understanding Evaluation) and SQuAD(Stanford Question Answering Dataset). In contrast, the time series analysis field lacks large-scale general benchmark datasets and large-scale challenges like those in image and language domains. The most famous time series forecasting competition is the M-Competitions, which started in 1982 and has only been held six times to date. The paper “Time Series Dataset Survey for Forecasting with Deep Learning” [68] highlights the lack of general benchmark datasets for TSF and analyzes the statistical characteristics of datasets across various domains, clustering them to provide easily accessible information for researchers. Recently, the development of foundation models for time series has begun, leading to the gradual emergence of general datasets for TSF. Details about TSF foundation models will be discussed in Section 4.3.
@@ 2.4 Evaluation Metrics
The evaluation metrics used in TSF play a crucial role in objectively comparing and assessing model performance. Common metrics for deterministic models include Mean Absolute Error (MAE) and Mean Squared Error (MSE), which are easy to use because they allow intuitive comparisons of prediction errors. On the other hand, probabilistic models that can reflect uncertainty typically use Continuous Ranked Probability Score (CRPS), which measures the accuracy of distribution forecasts by evaluating the difference between predicted and actual distributions. However, the limitations of these traditional evaluation metrics have been noted, leading to the proposal of various performance metrics.
@@
The distinct characteristics of time series data necessitate the use of diverse metrics in TSF. It is essential to assess not only the magnitude of errors but also how well the model learns the temporal patterns and various features of time series data. Table 4 categorizes evaluation metrics by type and explains each metric. Each metric assesses different characteristics of the model, helping to comprehensively understand model performance in various scenarios.
For instance, each data point in time series data has temporal dependencies, and it is necessary to evaluate how well the model learns these dependencies. Therefore, beyond simply measuring the magnitude of prediction errors with metrics like MAE or MSE, it is essential to evaluate how these errors change over time. Additionally, assessing relative errors is necessary to determine whether the model accurately captures seasonal or periodic patterns, making metrics like Mean Absolute Percentage Error (MAPE) or Symmetric Mean Absolute Percentage Error (sMAPE) necessary. Sometimes, metrics like R², which explains the variability of the entire dataset, are useful for examining long-term trends of increase or decrease in time series data. The continuous and irregular nature of time series data can mean large errors may occur at certain points, and metrics like Root Mean Squared Error (RMSE) are sensitive to large errors, allowing evaluation of the model’s ability to handle significant mistakes. To assess the tendency for prediction errors to increase over time, metrics such as Mean Forecast Error (MFE) or Cumulative Forecast Error (CFE) can be used to evaluate the accuracy of future predictions. Despite the availability of a wide range of metrics, MSE and MAE remain the most widely used due to their simplicity, intuitiveness, and widespread use in existing research, making them effective for objective performance comparisons.
@@
    Deterministic Error-Based Metrics Mean Absolute Error (MAE) $\frac{1}{n}\sum_{t=1}^{n}|y_{t}-\hat{y}_{t}|$ Measures the average magnitude of the errors in a set of predictions, without considering their direction.
@@
  Mean Absolute Percentage Error (MAPE) $\frac{1}{n}\sum_{t=1}^{n}\left|\frac{y_{t}-\hat{y}_{t}}{y_{t}}\right|\times 100$ Measures the size of the error in percentage terms.
@@
  Mean Forecast Error (MFE) $\frac{1}{n}\sum_{t=1}^{n}(y_{t}-\hat{y}_{t})$ Average of forecast errors, indicating bias.
@@
  Scaled Error Metrics Mean Absolute Scaled Error (MASE) $\frac{\frac{1}{n}\sum_{t=1}^{n}|y_{t}-\hat{y}_{t}|}{\frac{1}{n-1}\sum_{t=2}^{n}|y_{t}-y_{t-1}|}$ MAE scaled by the MAE of a naive forecast.

### v3

@@ Abstract
Time series forecasting is a critical task that provides key information for decision-making across various fields, such as economic planning, supply chain management, and medical diagnosis. After the use of traditional statistical methodologies and machine learning in the past, various fundamental deep learning architectures such as MLPs, CNNs, RNNs, and GNNs have been developed and applied to solve time series forecasting problems. However, the structural limitations caused by the inductive biases of each deep learning architecture constrained their performance. Transformer models, which excel at handling long-term dependencies, have become significant architectural components for time series forecasting. However, recent research has shown that alternatives such as simple linear layers can outperform Transformers. These findings have opened up new possibilities for using diverse architectures, ranging from fundamental deep learning models to emerging architectures and hybrid approaches. In this context of exploration into various models, the architectural modeling of time series forecasting has now entered a renaissance. This survey not only provides a historical context for time series forecasting but also offers comprehensive and timely analysis of the movement toward architectural diversification. By comparing and re-examining various deep learning models, we uncover new perspectives and presents the latest trends in time series forecasting, including the emergence of hybrid models, diffusion models, Mamba models, and foundation models. By focusing on the inherent characteristics of time series data, we also address open challenges that have gained attention in time series forecasting, such as channel dependency, distribution shift, causality, and feature extraction. This survey explores vital elements that can enhance forecasting performance through diverse approaches. These contributions help lower entry barriers for newcomers by providing a systematic understanding of the diverse research areas in time series forecasting (TSF), while offering seasoned researchers broader perspectives and new opportunities through in-depth exploration of TSF challenges.
@@ 1 Introduction
Time series forecasting (TSF) is a task that predicts future values based on sequential historical data [45]. It is utilized as a key decision-making tool in various fields, such as economics and finance, supply chain management, transportation, energy, weather, and healthcare [47, 1, 6, 179, 50, 231]. Such applications offer various opportunities, including cost reduction, increased efficiency, and enhanced competitiveness [47]. The diversity and complexity of time series data make forecasting challenging. In addition to apparent information, hidden patterns and irregular values pose challenges in learning temporal dependencies. [34]
In multivariate problems, additional factors such as channel correlation make the task even more difficult (Section 2.1). Furthermore, time series data exhibits different characteristics depending on the domain, and the various times and environments in which the series is collected result in significantly different patterns (Section 2.3). Because of this, TSF problems exhibit limited model generalizability, requiring diverse architectures and approaches. The increasingly complicated TSF problems are presenting researchers with growing challenges, which has recently led to the active development of new methodologies and algorithms to address these issues [141].
TSF is closely connected to core AI challenges, and solving these problems plays a crucial role in advancing the field of AI. For instance, the generalization of AI models is essential for ensuring stable performance on new data. However, time series data presents unique characteristics and domain-specific patterns that make generalization particularly challenging [94]. Solutions to generalization in TSF provide new insights that are applicable to other fields, such as Natural Language Processing (NLP) and Computer Vision (CV). The distribution of real-world time series data can change over time, and models must effectively handle these changes (Section 5.2). TSF plays a crucial role in addressing distribution shift issues, and the techniques developed to solve these problems can extend to other AI domains, such as image, video, and NLP. Moreover, attempts to understand models are crucial for improving their trustworthiness in real-world applications [93]. In mission-critical fields such as healthcare and finance, the explainability of time series prediction models is essential. For example, explaining how stock price predictions or patient vital sign forecasts are made can help users better understand and trust the model’s decisions. Thus, TSF research in explainability provides practical solutions for other AI applications (Section 5.6). Recently, TSF has evolved beyond simple forecasting, establishing itself as a practical platform for addressing key AI challenges. Solving these problems within TSF highlights the potential for AI to develop in a complementary and synergistic way.
@@
Fig. 2 shows the number of papers on time series forecasting presented at major AI and machine learning conferences. This explosive increase demonstrates the growing importance of TSF research in the AI and machine learning fields. As various studies addressing time series forecasting problems are being actively conducted, survey papers are also being frequently published. Over time, numerous survey papers have systematically organized the vast landscape of TSF, offering in-depth research that has provided valuable guidance and direction for researchers. However, existing survey papers still have room for improvement, particularly in addressing the inevitable increase in model diversity and the open challenges in the field, as summarized in Table 1. This survey aims to reorganize the extensive discourse on TSF by providing essential and comprehensive information. At the same time, it explores model diversity to deliver the latest TSF trends and uniquely highlights open challenges, clearly presenting them to readers.
@@
Models for TSF have undergone various stages of development over an extended period. In the past, statistical methods based on moving averages were predominantly used, which later evolved into traditional approaches such as Exponential Smoothing and ARIMA [15]. Machine learning techniques such as Tree Models [200] and Support Vector Machines (SVM) [42] have also been frequently used, but they had limitations in learning complex nonlinear patterns (Section 3.1). With the increase in available data and advancements in hardware computing power, various deep learning architectures such as MLPs [209], RNNs [100], CNNs [130], and GNNs [214] were developed, enabling the learning of more complex patterns. However, the performance of these early deep learning architectures was constrained by their intrinsic designs. To overcome these structural limitations, variants such as Long Short-Term Memory (LSTM) [98] and Temporal Convolutional Networks (TCN) [9] have been widely utilized (Section 3.2). Transformers [248], known for their ability to handle long-term dependencies, have demonstrated excellent performance in natural language processing and have been naturally extended to time series data as well. While Transformers have shown good performance in TSF and become widely popular, recent cases have shown that simple linear models can outperform Transformer models (Section 3.3). As a result, there has been a significant increase in reconsidering fundamental deep learning methodologies, along with growing interest in various architectures such as foundation models, diffusion models, and Mamba models (Section 4.2, 4.3, 4.4, 4.5). The Transformer model continues to improve in performance and still plays a significant role (Section 4.1). In this way, TSF has entered a renaissance of modeling, with various methodologies actively competing without being dominated by any single approach (Fig. 3). In this context, this survey offers two major strengths that set it apart from previous TSF survey papers.
@@
Second, we explore from the perspective of open challenges. Although numerous advanced architectures have resolved many issues, the core challenges in TSF continue to persist. In particular, issues such as channel correlation, distribution shifts, causality, and feature extraction (Section 5) remain significant challenges that need to be addressed (Fig. 4). This survey explores the latest methodologies aimed at addressing these challenges and provides readers with valuable insights for problem-solving. While some existing surveys address these TSF challenges, they often focus on only a few problems. In contrast, this paper takes an issue-driven approach, deeply examining seven key TSF problems and organizing new solutions.
@@
    Articles Focus Broad Architecture Review TSF Challenges Recent Work Reference
@@
    Forecast Methods for Time Series Data: A Survey · Categorization of various forecasting methods: statistical, ML, and DL
@@ A Review on Deep Sequential Models for Forecasting Time Series Data · Common deep sequential models
(Section 5.1)(Section 5.2)(Section 5.3)(Section 5.4)(Section 5.5)(Section 5.6)(Section 5.7)
@@
This survey covers the fundamental concepts of time series data and the problem definition of forecasting in Section 2, followed by an examination of the evolution of past methodologies in Section 3. Section 4 analyzes the key features of the latest models, and finally, Section 5, explores the open challenges in TSF and their solutions. Through this, readers will gain a broad understanding of the past and present of TSF research and acquire new ideas for future studies.
@@ 2.1 Time-Series Data
Time series data is a collection of sequential data points gathered at regular time intervals, representing a series of observations of phenomena that change over time. This temporal continuity allows for the understanding and analysis of phenomena that evolve according to time order. Each data point represents the state or value at a specific moment, and through the observation of these data points, various patterns such as long-term trends, seasonality, cyclicality, and irregularities can be recognized. These patterns provide valuable information for predicting future values or detecting changes at critical moments. By extracting and studying the meaningful information provided by time series data, researchers can develop practical applications to address challenges in various disciplines.
@@ 2.2 Time-Series Forecasting
Univariate forecasting refers to making predictions using only one variable. For example, predicting the next day’s temperature based solely on past temperature data from a weather station is a univariate forecast. The advantage of univariate forecasting is that the models are simple and computationally efficient. Since the data consists of a single variable, the model is relatively straightforward and easy to understand, making data collection and management easier. However, it may only utilize limited information as it cannot account for important external factors or interactions between different variables.
$\hat{x}_{t+1:t+h}=f({x}_{t-p:t})$ (1)
The formula representing UTSF is shown in Eq. (1). The model takes the past $p+1$ time steps, specifically the data from $t-p$ to $t$, as input to make predictions. Here, $\hat{x}_{t+1:t+h}$ represents the predicted values from $t+1$ to $t+h$. Single-step forecasting predicts only one value when $h=1$, while multi-step forecasting predicts multiple values when $h>1$.
@@
To incorporate other influencing variables, we can extend the model to a multivariate approach.
$\hat{\mathbf{X}}_{t+1:t+h}=f(\mathbf{X}_{t-p:t})$ (2)
The formula representing MTSF is shown in Eq. (2). $\mathbf{X}_{t-p:t}$ is a set of vectors over the past $p+1$ time steps, where each vector contains the values of multiple variables at time $t$. This model takes multiple variables from the past $p+1$ time steps, specifically $\mathbf{X}_{t-p:t}$, as input to generate the multivariate predicted values from $t+1$ to $t+h$, denoted as $\hat{\mathbf{X}}_{t+1:t+h}$. This multivariate approach allows for more comprehensive modeling by considering both the target variable’s history and the effects of other relevant factors. Fig. 7 illustrates how a multivariate forecasting model uses a past lookback window to predict future intervals.
@@ 2.3 Time Series Forecasting Datasets
Time series data exhibit a greater variety of domain-specific characteristics compared to Natural Language Processing (NLP). These differences in dataset characteristics make time series forecasting problems more diverse and complex, reducing the generalizability of time series forecasting models. Consequently, developing foundation models for TSF is more challenging compared to NLP. Time series data lack the well-designed vocabulary and grammar found in NLP, and obtaining vast amounts of data is more difficult than in fields like Computer Vision (CV) or NLP. In the CV field, there are general benchmark datasets such as MNIST and ImageNet, while in NLP, there are datasets like GLUE (General Language Understanding Evaluation) and SQuAD (Stanford Question Answering Dataset). In contrast, the time series analysis field lacks large-scale general benchmark datasets and large-scale challenges like those in image and language domains. The most famous time series forecasting competition is the M-Competitions, which started in 1982 and has only been held six times to date. The paper “Time Series Dataset Survey for Forecasting with Deep Learning” [91] highlights the lack of general benchmark datasets for TSF and analyzes the statistical characteristics of datasets across various domains, clustering them to provide easily accessible information for researchers. We do not delve into the characteristics of the datasets in detail, but instead provide a brief overview of the key datasets. Recently, the development of foundation models for time series has begun, leading to the gradual emergence of general datasets for TSF. Details about TSF foundation models will be discussed in Section 4.3.
@@ 2.4 Evaluation Metrics
The evaluation metrics used in TSF play a crucial role in objectively comparing and assessing model performance. Common metrics for deterministic models include Mean Absolute Error (MAE) and Mean Squared Error (MSE), which are easy to use because they allow intuitive comparisons of prediction errors. On the other hand, probabilistic models that can reflect uncertainty typically use Continuous Ranked Probability Score (CRPS), which measures the accuracy of distribution forecasts by evaluating the difference between predicted and actual distributions [168]. However, the limitations of these traditional evaluation metrics have been noted, leading to the proposal of various performance metrics [242].
@@
The distinct characteristics of time series data necessitate the use of diverse metrics in TSF. It is essential to assess not only the magnitude of errors but also how well the model learns the temporal patterns and various features of time series data. Table 4 categorizes evaluation metrics by type and explains each metric.
Error-based metrics for deterministic models measure the difference between predicted and actual values. Explained variance metrics assess how well the model explains the variance of the data [39]. These metrics evaluate how well the model captures key patterns such as seasonality or trends in the data, thereby aiding in the structural understanding of the predictions. Error metrics focus on accuracy, but they do not assess overfitting, whereas explained variance evaluates the model’s generalization ability. Model selection metrics help identify the optimal model [196]. These metrics take into account factors such as sample size and the number of parameters, playing a key role in selecting the optimal model. Thus, they are particularly useful for choosing the most appropriate model for a given dataset. For instance, in noisy environments, MAE is preferred over MSE due to its reduced sensitivity to outliers. Additionally, when evaluating non-stationary data, it is important to use metrics that effectively capture changes in trends and seasonality. In such cases, cumulative error-based metrics or explained variance metrics can be useful.
Probabilistic model evaluation metrics, on the other hand, account for uncertainty, enabling the assessment of both the confidence and variability of predictions. These metrics help evaluate not only the predicted values but also the uncertainty surrounding them. Beyond basic error-based metrics, several advanced metrics are used. Interval Metrics assess the confidence intervals of the predictions, considering both the predicted values and their associated uncertainty [119]. These metrics are especially useful in fields such as financial market forecasting and risk management. Quantile Metrics evaluate the accuracy of specific quantiles of predicted values, making them particularly useful when assessing how well various quantiles match the actual values in situations with high uncertainty [126]. Quantile Metrics are beneficial in areas such as weather forecasting and insurance, where forecasts are assessed across different risk levels. Sharpness Metrics, which evaluate the width of predicted confidence intervals, are used to assess how narrow and accurate the prediction intervals are [308]. This metric is valuable in logistics and supply chain forecasting, where minimizing the width of prediction intervals is crucial.
Thus, various evaluation metrics are crucial in assessing model performance, and each should be chosen based on the specific characteristics of the data and the requirements of the problem. In real-world scenarios, the appropriate use of metrics allows for the selection of the most effective model, which can then be leveraged for efficient decision-making. In contrast, in the field of deep learning research, MSE and MAE remain widely used due to their simplicity, intuitiveness, and the fact that they have been extensively utilized in previous studies, making them suitable for performance comparison [35].
@@
    Deterministic Error-Based: Absolute Mean Absolute Error (MAE) $\frac{1}{n}\sum_{t=1}^{n}|y_{t}-\hat{y}_{t}|$ Measures the average magnitude of the errors in a set of predictions, without considering their direction.
@@
  Error-Based: Relative Mean Absolute Percentage Error (MAPE) $\frac{1}{n}\sum_{t=1}^{n}\left|\frac{y_{t}-\hat{y}_{t}}{y_{t}}\right|\times 100$ Measures the size of the error in percentage terms.
@@
  Error-Based: Cumulative Mean Forecast Error (MFE) $\frac{1}{n}\sum_{t=1}^{n}(y_{t}-\hat{y}_{t})$ Average of forecast errors, indicating bias.
@@
  Error-Based: Scaled Mean Absolute Scaled Error (MASE) $\frac{\frac{1}{n}\sum_{t=1}^{n}|y_{t}-\hat{y}_{t}|}{\frac{1}{n-1}\sum_{t=2}^{n}|y_{t}-y_{t-1}|}$ MAE scaled by the MAE of a naive forecast.